A 200g Puck is moving at 5m/s on an ice rink. It skids off the ice unto a floor, where the coefficient of kinetic friction between it and the floor is 0.25. How far does it move along the floor before stopping?
1
Expert's answer
2020-06-30T18:24:09-0400
"v^2=2\mu gs\\5^2=2(0.25)(9.8)s\\s=5.1\ m"